OpenStudy (jamesreed30):

Express the required measurement in exact form what is the radius of a circle in which a 30° arc is two pi inches long

OpenStudy (wonderwomen):

Okay I hope this is the right answer but let me explain to you first number of arcs in a full circle: 360 / 30 = 12 circumference of that circle is hence: 12 * 2pi = 24pi 2*pi*r = 24pi r = 24pi / 2pi = 12 inches

OpenStudy (wonderwomen):

Hope this helps :)

OpenStudy (jamesreed30):

So 12 would be the answer?